# Investigating LLM Capabilities on Long Context Comprehension for Medical Question Answering

### v1

@@ 4.1 Datasets
Context Sources: We leverage MIMIC-III (Johnson et al., 2016) and MIMIC-IV (Johnson et al., 2023) comprising de-identified EHRs covering hospital admissions and ICU stays. They support single- and multi-note settings and long contexts. Our data includes: clinical notes, discharge summaries and radiology reports per patient over time.
QA Datasets: We test on four EHR QA benchmarks spanning different tasks given MIMIC-based underlying context. We only use QA pairs curated or validated by clinical experts to ensure medical validity. We provide a brief dataset overview below, with specifics summarized in Table 1.
CliniQG4QA (Yue et al., 2020): An extractive span benchmark originally created for domain adaptation purposes based on MIMIC-III clinical notes. While the 8,824-QA pair dev set is machine-generated (MG), the 1,287 QA test pairs are expert-annotated using the MG questions for reference. Answers were either generated or verified by three clinical experts. Here we only use the expert-annotated test set.
RadQA (Soni et al., 2022): An extractive QA dataset comprising 3,074 physician-crafted questions and 6,148 answer spans from the Findings and Impressions sections of radiology reports in MIMIC-III. Many questions correspond to multiple sentences and are shown to require different types of reasoning to answer. Annotation is carried out by two human annotators with medical understanding. Here we use the train set and filter out Unanswerable QA pairs resulting in 3,509 pairs.
EHR-DS-QA (Kotschenreuther, 2024): A large synthetic QA corpus of 156,599 pairs generated by two Llama-based (Touvron et al., 2023) LLMs and guided by predefined prompt templates on discharge summaries and notes from MIMIC-IV. A subset of 506 pairs were physician-verified of which we retain 478, marked as correct.
EHRNoteQA (Kweon et al., 2024): Built on MIMIC-IV discharge summaries, it contains 962 QA pairs authored by GPT-4 and iteratively refined by three clinicians. Questions span a diverse set of topics, each corresponding to multiple discharge summaries ($\sim$2.3 per patient). The dataset supports both open-ended and MC formats.
@@ 4.2 Context Formulation
Since most of the datasets are based on a single note/report (CliniQG4QA, RadQA, EHR-DS-QA) or up to three discharge summaries (EHRNoteQA), we identify a large amount of longitudinal patient supplementary content that remains unused. For a more realistic clinical setting, we leverage our context sources, MIMIC-III and MIMIC-IV, to augment few-note context with each patient’s longitudinal notes. We then investigate how LLMs leverage long-form context to answer patient-specific questions. We believe that this setting, (Include All), presents a more realistic scenario where a model would have no prior knowledge of the relevant document and would therefore need to process the entire patient note history [^3] MIMIC preprocessing is described in Appendix C.2..
We explore four data inclusion scenarios depicted in Fig. 1, allowing comprehensive analysis of model sensitivity to context size and noise, ability to handle supplementary information and memorization:
@@
    Exclude All: Exclusion of all clinical notes - assessing model note memorization.
@@
    Exclude Relevant: Exclusion of relevant clinical notes - assessing usefulness of supplementary material.
@@
    Include All: See above.
@@
    Include Related: Inclusion of only dataset relevant clinical note types: discharge summary(ies), note or radiology report (as marked in each underlying dataset) - assessing model’s ability to leverage relevant context information for patient-centric QA.
Context Segmentation: For the Include All setting we report context-length-based performance across four bins: i) Short context: 0–8K, ii) Medium context: 8–16K, iii) Large context: 16–32K, iv) Extended context: 32–128K tokens.
@@ 4.3 Models and Experimental Setup
LLMs: To explore the effect of LLM model size, domain specialization and reasoning capabilities in a controlled way we study four open Qwen2.5 models (Team and others, 2024), namely:
@@
  • HuatuoGPT-o1-Qwen-7B (Chen et al., 2024): medical LLM based on Qwen2.5-7B-Instruct
@@
Retrieval Settings: We implemented a hybrid retrieval pipeline designed to properly handle the semantic, lexical, and query complexity of medical QA. Our pipeline involves a combination of three different retrievers (more details are in §3.3):
  • Dense Retrieval based on the leading open embedding model Qwen3-Embedding-8B (Zhang et al., 2025).
  • Sparse Retrieval performed using BM25 selected after ablating two commonly used sparse retrievers (Table 8 in Appendix).
  • Late Interaction Reranking with Reason-ModernColBERT, a ColBERT (Khattab and Zaharia, 2020) model finetuned on the ReasonIR dataset (Shao et al., 2025) leading to very high performance on reasoning-intensive retrieval benchmark (Su et al., 2024). Due to its performance intensive nature, we use it as a final reranker in our pipeline.
Experimental Setup: We defer inference specifics to Appendix A and prompt formulation details in Appendix C.3.
@@ 5 Results and Discussion
Full context performance for the different settings for all datasets and models is shown in Table 2. For RAG, Table 3 shows different evaluated RAG settings across 3 selected models on the 2 MIMIC-IV based datasets. We excluded QwQ:32B due to time and budget constraints. Similarly, in Table  4 we only evaluate 7B sized models using a single RAG strategy, due to the task nature and limited context size of the MIMIC-III based datasets.
Model Size, Context Size and Performance
Consistent with general findings, larger models consistently perform better across all tasks (Table 2). Qwen2.5:32b and QwQ:32b generally outperformed the 7b parameter models across most metrics and tasks. The performance gap was particularly pronounced in tasks requiring reasoning or synthesizing of information from multiple sources, namely EHRNoteQA and RadQA. However, smaller models performed better in settings excluding relevant information or completely removing the context. This suggests that they are more prone to memorization.
The ‘Include Related’ setting outperforms ‘Include All’ on most datasets and metrics except for the MC task where no benefits from filtering are seen for 3 out of the 4 models. Fig. 3 shows results for Qwen2.5:32B-Instruct over different context sizes. We observed similar trends between other models and context size (see Appendix, Fig.  LABEL:fig:model_performance_context). Worth noting that for context range (8-16K) EHR-DS-QA exhibits a dip in performance across all metrics and models, attributable to noise (see §5.1).
@@
Reasoning and Fine-tuning
The impact of reasoning-focused models and medical fine-tuning is mixed. The QwQ:32b general reasoning model showed improved results on Table 2 over other models overall on our two reasoning datasets (RadQA, EHRNoteQA). This suggests that specialized reasoning models can offer an advantage in tasks that demand complex inference across multiple sources of information, but this benefit may be more apparent in larger model sizes where reasoning capabilities can be effectively leveraged without compromising other essential skills.
By contrast, the medically fine-tuned reasoning model, HuatuoGPT-o1-Qwen2.5:7b, did not show any improvement over its standard instruction-tuned base model Qwen2.5:7b-instruct, in line with studies revealing that biomedical LLMs often lead to reduced performance (Dorfner et al., 2024; Dada et al., 2024).
@@
RAG Performance
Table 3 provides insights into the effectiveness of different RAG strategies per task and dataset. RAG demonstrated clear performance improvement for the single-note generative task of EHR-DS-QA. In this scenario, the chunk-in-context strategy consistently provided the best results. For the more complex, multi-note reasoning task of EHRNoteQA, the hierarchical RAG strategy was the best with few exceptions. The results also show similar patterns on MC indicating clear performance improvement for all models on the RAG HIR 7 setting. These findings are consistent across different context sizes (Fig.  4).
@@
Worth noting that using RAG, the medically fine-tuned reasoning model HuatuoGPT-o1-Qwen2.5:7b was able to score the best results on multiple metrics across the 2 datasets. This shows that RAG could benefit smaller models that struggle in FC settings.
While the extractive datasets have shorter contexts overall, Table 4 shows that the RAG setting yields the best performance.
@@
Metrics Insights
Fig. 5 provides insights into the relation between models, metrics, and datasets while focusing on the open-ended generative tasks for the ‘Include All’ and ‘Include Related’ settings combined. Instruct models lead on single-note non-reasoning tasks (EHR-DS-QA). Larger models lead on EHRNoteQA: instruct models are better based on semantic and NLI metrics, while reasoning ones score better on LLM as a Judge metrics  [^4] We include all-context and long-context correlations between metrics in Appendix LABEL:sec:appendix.metric_correlation.
FC vs RAG in QA Tasks
We did a qualitative analysis on the performance of FC and RAG on the generative task across both datasets detailed in Appendix LABEL:sec:appendix.app-rag-vs-fc with examples.
We provide a summary of the main points below:
  • FC performs better for questions requiring comprehensive understanding of the entire document or when answers are located in structured sections, such as summaries, lists, or temporal sequences.
  • RAG outperforms FC for questions that require retrieving specific details or synthesizing information scattered across multiple parts of the document.
  • RAG tends to perform better overall in datasets with complex or lengthy documents, likely due to its ability to focus on relevant information and reduce noise.
  • There is mixed evidence suggesting FC might perform better for tasks involving inferential reasoning or identifying the absence of information.
@@
5.1 Error Analysis
We analyze failure instances that complement the quantitative trends and provide diagnostic guidance for practical deployment. We summarize three recurrent themes and refer to Appendix LABEL:sec:appendix.case_studies for full cases and diagnostic tables.
Data noise. We observe inconsistencies and ambiguous gold answers in some pairs across all datasets (more profoundly in EHR-DS-QA) that correlate with strange dips in performance at mid-range context range windows. Fan et al. (2024) also observed similar performance dips.
Numerical Values. Questions requiring numeric values (e.g., vitals, dosages, dates) are sensitive to formatting, rounding, or range expressions.
Metric Insights. We did a qualitative analysis to understand the reasons behind metrics’ disagreements across six metrics on the two generative datasets (see details in Appendix LABEL:app:appendix.metric_insights).
We found 233 disagreements in EHRNoteQA and 47 in EHR-DS-QA. These are mainly due to:
  1. 1.
    Correct/Complete but Unfaithful (addition of unsupported details). Most common in EHRNoteQA (173/233; 74%) and present in EHR-DS-QA (17/47; 36%).
  2. 2.
    Correct answer but low surface overlap (lower Meteor/Bert) appears in both datasets (EHRNoteQA: 20%; EHR-DS-QA: 26%).
  3. 3.
    Correct answer but low NLI entailment occurs due to negation/clinical phrasing (EHRNoteQA: 17%; EHR-DS-QA: 21%).
  4. 4.
    EHR-DS-QA shows several short, risk/negation items where NLI flags contradiction despite clinically aligned answers.
This points to the need for thinking about more reasoning appropriate metrics.
@@ 6 Conclusion
We studied long-context, patient-centric clinical QA across EHR-grounded datasets, contrasting Full-Context prompting with Retrieval-augmented generation (RAG). Our analysis shows that carefully scoped context (e.g., Include DS) often outperforms feeding all notes, and that hybrid RAG pipelines with reranking are competitive, especially for specific factual queries and multi-note synthesis. Larger models generally help, but performance remains sensitive to context length, evidence placement, and task formulation. Future directions include: (i) scaling beyond single-note or single-source assumptions toward richer multi-note, multi-visit reasoning; (ii) advancing temporal and causal reasoning over longitudinal records, with explicit timeline grounding; (iii) strengthening evaluation via clinically faithful, judge-robust rubrics and cross-metric reliability analyses.
@@
Despite the in depth study of literature and our initial ablations in selecting a competitive RAG methodology, our work does not exhaustively examine the full potential of RAG under different settings, i.e. different retrievers and chunk sizes, in long-form medical QA. Finally, while we focus on Qwen-based models due to their competitive performance, we acknowledge that an evaluation of a wider range of LLMs could offer a more complete picture of the LLM landscape in long-form medical QA. Our analysis has also pointed to the limitations of current evaluation metrics, both in terms of assessing reasoning as well as lack of transparency in LLM as a judge metrics and how faithfulness can be truly assessed currently when additional information, not present in the benchmark, is added.
@@ C.3 Prompt Formulations
We used a simple one shot prompt structure and tailored it explicitly for each task formulation:

### v2

@@ 4.1 Datasets
We leverage MIMIC-III (Johnson et al., 2016) and MIMIC-IV (Johnson et al., 2023), comprising de-identified EHRs covering hospital admissions and ICU stays, as the underlying context. They support single- and multi-note settings and include: clinical notes (CN), discharge summaries (DS) and radiology reports (RR) per patient over time.
Our four selected QA datasets only use QA pairs curated or validated by clinical experts to ensure medical validity. We provide a brief dataset overview below, also summarized in Table 1:
CliniQG4QA (Yue et al., 2021): An extractive span benchmark based on MIMIC-III. While most of its pairs are machine-generated (MG), the 1,287 QA test pairs are either generated or verified by three clinical experts using the MG questions for reference. We only use the expert-annotated set.
RadQA (Soni et al., 2022): An extractive dataset comprising 3,074 physician-crafted questions and 6,148 answer spans from the Findings and Impressions sections of RR in MIMIC-III. Many questions correspond to multiple sentences and require different types of reasoning. Annotation is carried out by two human annotators. After filtering out for Unanswerable pairs we obtain 3,509 pairs.
EHR-DS-QA (Kotschenreuther, 2024): A synthetic QA corpus of 156,599 pairs generated by two LLMs and guided by predefined prompt templates on DS and CN from MIMIC-IV. A subset of 506 pairs were physician-verified of which we retain 478, marked as correct.
EHRNoteQA (Kweon et al., 2024): Built on MIMIC-IV, it contains 962 QA pairs authored by GPT-4 and iteratively refined by three clinicians. Questions span a diverse set of topics, each corresponding to multiple DS ($\sim$2.3 per patient), while supporting both open-ended and MCQA formats.
@@ 4.2 Context Formulation
While most of the datasets are based on a single note/report (CliniQG4QA, RadQA, EHR-DS-QA) or up to three DS (EHRNoteQA), there is a large amount of longitudinal patient supplementary content that remains unused. Aiming for a realistic clinical setting where there is no a-priori knowledge of the most relevant note, we leverage the MIMIC resources to augment the context with each patient’s longitudinal notes [^2] MIMIC preprocessing is described in Appendix C.2., resulting in long context from the entire patient’s note history (Include All). We then investigate how LLMs leverage long-form context to answer patient-specific questions.
We explore four data inclusion scenarios depicted in Fig. 1, allowing comprehensive analysis:
@@
    Exclude All: Exclusion of all notes - assessing model note memorization.
@@
    Exclude Relevant: Exclusion of relevant notes - assessing usefulness of supplementary material.
@@
    Include All: Inclusion of all patient notes - assessing LC processing of varying importance.
@@
    Include Related: Inclusion of only the most relevant note(s) (DS, CN or RR) as marked in each dataset - assessing model’s ability to leverage the most relevant context information.
For the Include All setting we report context-length-based performance across four token bins: Short context: 0–8K, Medium context: 8–16K, Large context: 16–32K, Extended context: 32–128K.
@@ 4.3 Models and Experimental Setup
LLMs: To explore the effect of model size, domain specialization and reasoning capabilities we study four open Qwen2.5 (Team et al., 2024b) single-family models, ensuring a controlled comparison:
@@
  • HuatuoGPT-o1-Qwen-7B (Chen et al., 2024b): medical LLM based on Qwen2.5-7B-Instruct
@@
Evaluation: Through comparisons against Selene-8B (Alexandru et al., 2025) and Prometheus-8x7B v2.0 (Kim et al., 2024), we finally selected Qwen-2.5-32B-Instruct for the LLM-as-a-judge for stability and agreement (see Appendix LABEL:sec:judgecomps).
Retrieval Settings: Based on biomedical QA retrieval research that demonstrates the superiority of combining sparse and dense retrievers (see §3.2), we designed a hybrid retrieval approach to properly handle the semantic, lexical, and query complexity. The selected model components are enlisted below:
  • Dense Retrieval using the open embedding model Qwen3-Embedding-8B (Zhang et al., 2025).
  • Sparse Retrieval using BM25, selected after ablating two sparse retrievers (see Table 9).
  • Late Interaction Reranking with Reason-ModernColBERT, a ColBERT (Khattab and Zaharia, 2020) model finetuned on the ReasonIR dataset (Shao et al., 2025) demonstrating high performance on a reasoning-intensive retrieval benchmark (Su et al., 2025).
We explored two chunk inclusion strategies: i) direct chunk inclusion (RAG), ii) hierarchical parent note inclusion (RAG HIR).
Experimental Setup: LLM inference and prompt formulation specifics are in Appendices A and C.3.
@@ 5 Results and Discussion
5.1 Main Results
Performance for the different inclusion settings across datasets and models is shown in Table 2.
Model Size, Tasks and Inclusion Settings:
Consistent with general findings, larger models perform better across all tasks (Table 2). Qwen2.5:32B and QwQ:32B generally outperform the 7B parameter models across most metrics and tasks. The performance gap is particularly pronounced in tasks requiring reasoning or information synthesis from multiple sources, namely EHRNoteQA and RadQA. However, smaller models perform better in settings excluding relevant information or completely removing the context. This suggests that they are more prone to memorization. Additionally, unlike the rest of the tasks, MC shows strong evidence of memorization with a performance of 59.24% in the ‘Exclude All’ setting. While MC demonstrates 90%+ performance, open-ended and extractive tasks remain challenging for LLMs.
The ‘Include Related’ setting outperforms ‘Include All’ on most datasets and metrics except for MC. Despite that, the ‘Exclude Related’ setting outperforms the ‘Exclude All’ by a margin across the board demonstrating the usefulness of the supplementary patient material and suggesting that while the ‘Include All’ setting includes useful additional information, LLMs struggle to process long context and to identify the most relevant information.
@@ 5.1 Main Results
Reasoning and Fine-tuning:
The impact of reasoning-focused models and medical fine-tuning is mixed. The QwQ:32B general reasoning model showed improved results on Table 2 over other models overall on RadQA, EHRNoteQA - the reasoning datasets. This suggests that reasoning models can offer an advantage in tasks that demand complex inference across multiple sources of information, but this benefit may be more apparent in larger model sizes where reasoning capabilities can be effectively leveraged without compromising other essential skills. By contrast, the medically fine-tuned reasoning model, HuatuoGPT-o1-Qwen2.5:7B, did not show any improvement over its standard instruction-tuned base model Qwen2.5:7B-instruct, in line with studies revealing that biomedical LLMs often lead to reduced performance (Dorfner et al., 2024; Dada et al., 2025).
@@
Context Size:
Fig. 3 shows results for Qwen2.5:32B-Instruct over different context sizes. Generally, performance decreases across metrics with increased context size on open-ended QA particularly when multi-note reasoning is required (EHRNoteQA), showing that even large LLMs are challenged by long context, with similar trends observed across models (see Appendix, Fig.  LABEL:fig:model_performance_context). Worth noting that for context range (8-16K) EHR-DS-QA exhibits a dip in performance across all metrics and models. This is likely attributed to data noise, based on manually observed inconsistencies and ambiguous gold answers in some pairs especially in EHR-DS-QA, correlating with strange dips in performance at mid-range context windows. Fan et al. (2025a) and Ma et al. (2025) observed similar dips.
@@
RAG Performance:
Table 3 shows the two evaluated chunk settings on the open-ended generative datasets [^3] We excluded QwQ:32B due to time and budget constraints. RAG demonstrated clear performance improvement for the single-note generative task of EHR-DS-QA. In this scenario, the direct chunk inclusion strategy consistently provided the best results. For the more complex, multi-note reasoning task of EHRNoteQA, the hierarchical RAG strategy (RAG HIR) was the best with few exceptions. The results also show similar patterns on MC indicating clear performance improvement for all models on the RAG HIR 7 setting. Our findings are consistent across medium and larger (8K+) context sizes (Fig.  4). Worth noting that using RAG, the medically fine-tuned reasoning model HuatuoGPT-o1-Qwen2.5:7B scored the best results on multiple metrics across the open-ended datasets, showing that RAG could benefit even more so the smaller models that struggle in Full-context (FC) settings (‘Include All’).
In Table  4 we only evaluate 7B sized models using a single RAG setting, due to the task nature and limited context size of the extractive datasets. While the extractive datasets have shorter contexts overall RAG still yields the best performance.
Category Total Favored by RAG Favored by FC SEE 18 12 (66.7%) 6 (33.3%) TLR 10 3 (30.0%) 7 (70.0%) CNS 10 6 (60.0%) 4 (40.0%) ANI 3 2 (66.7%) 1 (33.3%) Total 41 23 (56.1%) 18 (43.9%)
Table 5: Breakdown of sampled favored by RAG vs FC across four categories in open-ended QA.
5.2 Quantitative Analysis: RAG vs FC
In our analysis we gathered challenging cases of disagreement between metrics (LLM Correctness, NLI Entailment, and Bio-F1) across EHRNoteQA and EHR-DS-QA. Metrics were first z-score standardized and samples were gathered for cases were some metrics belonged to the top 60th percentile while another metric fell within the bottom 40th percentile. The 41 examples gathered were then manually examined in order to categorize them across QA types and determine which answer between RAG and FC is more reliable. The rubric is provided below:
  • Specific Entity Extraction (SEE): Requiring retrieval of high cardinality facts such as medications and lab values or surgical procedures.
  • Temporal & Longitudinal Reasoning (TLR): Requiring tracking of changes over time, chronological event ordering or visit comparisons.
  • Clinical Nuance & Status (CNS): Questions regarding discharge instructions, patient mental status, or synthesized clinical course summaries.
  • Absence or Negative Information (ANI): Requiring the identification of absent information.
@@ 5.2 Quantitative Analysis: RAG vs FC
In Table 5, RAG shows higher performance across the SEE, CNS and ANI categories, while FC leads in the TLR one, demonstrating that while RAG effectively preserves the relevant information for challenging cases, it falls behind when temporal reasoning is required. A further qualitative analysis providing insights on cases where RAG or FC perform better (see Appendix LABEL:sec:appendix.app-rag-vs-fc) showcases findings on par with the quantitative ones, with the exception of ANI where FC has more promising although inconsistent performance.
5.3 Metrics and Insights
Metric Comparisons: Fig. 5 provides insights into the relation between models, metrics, and datasets while focusing on the open-ended generative tasks for the ‘Include All’ and ‘Include Related’ settings combined. Instruct models lead on the single-note non-reasoning task (EHR-DS-QA), while larger models lead on the multi-note reasoning one (EHRNoteQA). Instruct models are better based on semantic and NLI metrics, while reasoning ones score better on LLM-as-a-judge metrics [^4] We include all-context and long-context correlations between metrics in Appendix LABEL:sec:appendix.metric_correlation.
Metric Insights: We performed a qualitative analysis to investigate the reasons behind metrics’ disagreements on our open-ended datasets (see Appendix LABEL:app:appendix.metric_insights). The 233 disagreements in EHRNoteQA and 47 in EHR-DS-QA are mainly due to:
  • High Correctness/Completeness but low Faithfulness in predictions due to the addition of unsupported details - in both datasets but more pronounced in EHR-DS-QA.
  • High Correctness but with low surface overlap between the gold and the prediction (lower METEOR/BERTScore) - in both datasets.
  • High Correctness but with either low NLI entailment or high NLI contradiction occurring due to either existence of negation or shortness in prediction or differences in clinical phrasing - in both datasets.
These point to the need for thinking about more reasoning appropriate metrics [^5] Manually analyzed failure cases are in Appendix LABEL:sec:appendix.case_studies..
@@ 6 Conclusion
We studied long-context, patient-centric clinical QA across EHR-grounded datasets, contrasting Full-Context (FC) prompting with Retrieval-Augmented generation (RAG). Our work shows that highly-relevant context often outperforms feeding all notes, likely due to LLMs struggling with long context. Hybrid RAG pipelines with reranking overperform FC, especially for specific factual queries and multi-note synthesis. Larger models generally help especially for reasoning datasets, but performance remains sensitive to context length and task formulation, while medically fine-tuned models fall behind. Future directions include: (i) scaling beyond single-note or single-source assumptions toward richer multi-note, multi-visit reasoning; (ii) advancing temporal and causal reasoning over longitudinal records, with explicit timeline grounding; (iii) strengthening evaluation via clinically faithful, judge-robust rubrics and cross-metric reliability analyses.
@@
Despite the in-depth study of literature and our initial ablations in selecting a competitive RAG methodology, our work does not exhaustively examine the full potential of RAG under different settings, i.e. different retrievers and chunk sizes, in long-form medical QA. Finally, while we focus on Qwen-based models due to their competitive performance, we acknowledge that an evaluation of a wider range of LLMs could offer a more complete picture of the LLM landscape in long-form medical QA. Our analysis also highlights limitations in current evaluation metrics, specifically regarding the assessment of reasoning, the lack of transparency in LLM-as-a-judge frameworks, and the difficulty of accurately measuring faithfulness when benchmarks are supplemented with additional information.
@@ C.3 Prompt Formulations
We used a simple zero-shot prompt structure and tailored it explicitly for each task formulation: